CK Company stockholders expect to receive a year-end dividend of $5 per share and then immediately sell their shares for $115 dollars per share. If the required rate of return for the stock is 20 percent, what is the current value of the stock?
Answer #1

Solution:

The formula for calculating the required rate of return for a stock is

Required rate of return = ( Selling price + Dividend per share – Current Purchase price ) / Current Purchase price

As per the information given in the question we have

Required rate of return = 20 % = 0.20 ; Selling price = $ 115 per share ; Dividend per share = $ 5 ; Current Purchase price of stock = Current value of stock = To find

Let the current value of stock = ‘ x ‘

Applying the above information in the formula we have

0.20 = ( $ 115 + $ 5 – x ) / x

0.20x = $ 120 – x

0.20x + x = 120

1.2x = 120

x = 120 / 1.2 = 100

Thus the current value of stock = $ 100